OpenStudy (sherribelle):

I'm doing Arithmetic Sequence and I need help. The first term of an arithmetic sequence is 42. The rule an = an-1 + 8 can be used to find the next term of the sequence. Explain how to write the explicit rule for the arithmetic sequence from the given information. I just need help setting it up.

OpenStudy (sherribelle):

Wait wait, wouldn't you just add 8? Or would you subtract it?

OpenStudy (dumbcow):

you would add 8

OpenStudy (dumbcow):

explicit formula \[a_n = a_1 + d(n-1)\] a1 = 42 d = 8

OpenStudy (sherribelle):

So an-1 = 42? So it would look like an = 42 + 8

OpenStudy (sherribelle):

Well really an = 42 (n - 1)8